### v1

# rowhighlight linewidth=1pt, linecolor=black, innertopmargin=4pt, innerbottommargin=4pt, innerleftmargin=4pt, innerrightmargin=4pt, backgroundcolor=lightgray!15

Abstract
Sparse Autoencoders (SAEs) have demonstrated significant promise in interpreting the hidden states of language models by decomposing them into interpretable latent directions. However, training SAEs at scale remains challenging, especially when large dictionary sizes are used. While decoders can leverage sparse-aware kernels for efficiency, encoders still require computationally intensive linear operations with large output dimensions. To address this, we propose KronSAE, a novel architecture that factorizes the latent representation via Kronecker product decomposition, drastically reducing memory and computational overhead. Furthermore, we introduce mAND, a differentiable activation function approximating the binary AND operation, which improves interpretability and performance in our factorized framework.
1 Introduction
Interpreting dense embeddings from large language models remains a central challenge for transparency and controllability in AI systems (Elhage et al., 2023; Heap et al., 2025). Sparse autoencoders (SAEs) have emerged as powerful tools for uncovering human-interpretable features within neural activations by enforcing activation sparsity to induce discrete-style dictionaries (Elhage et al., 2023; Gao et al., 2025; Cunningham et al., 2024). These dictionaries facilitate rigorous circuit-level semantic analysis and concept discovery, enabling fine-grained probing of model internals (Elhage et al., 2023; Cunningham et al., 2024).
However, naively scaling SAEs to the widths demanded by modern transformers leads to prohibitive compute costs—both in memory and per-step FLOPs—limiting their applicability to frontier models. Gated SAEs (Rajamanoharan et al., 2024a) address this by learning continuous sparsity masks via lightweight gating networks, achieving a Pareto improvement on the reconstruction–sparsity trade-off. Switch SAEs (Mudide et al., 2025) leverage conditional computation by routing activations among smaller expert SAEs, reducing computation by activating only a subset of experts per input. In a work Gao et al. (2025) an acceleration of TopK SAE was proposed that utilizes an optimized kernel based on efficient sparse–dense matrix multiplication. Despite these advances, the encoder remains largely unoptimized: it still performs a dense projection into the full dictionary, incurring high computational cost and limiting scalability.
In this paper, we introduce KronSAE, a novel SAE encoder that directly tackles this bottleneck. By decomposing the latent space into head-wise Kronecker factors and embedding a smooth, differentiable AND-like gating mechanism, KronSAE reduces both parameters and compute overhead while preserving reconstruction fidelity.
This work makes three primary contributions:
  1. 1.
    We identify the encoder projection as the principal scalability bottleneck in sparse autoencoders, conducting a FLOPs-aware analysis of its impact and demonstrating that targeted encoder optimizations can significantly improve computational performance while maintaining reconstruction quality.
  2. 2.
    We propose KronSAE, a Kronecker-factorised sparse autoencoder equipped with the novel $\operatorname{mAND}$ activation function. The design significantly reduces encoder cost and is compatible with existing sparse decoder kernels.
  3. 3.
    We show on multiple language models that KronSAE improves reconstruction fidelity, decreases feature absorption, and yields more interpretable latents under fixed compute.
2 Related Work
Sparse Autoencoders.
Sparse autoencoders (SAEs) enforce sparsity constraint on a latent code. Early work demonstrated that SAEs can uncover human-interpretable directions in deep models (Cunningham et al., 2024; Templeton et al., 2024), but extending them to the large latent sizes ($F\gg d$) for modern language models with large hidden dimension $d$ is expensive: each forward pass still requires a dense $\mathcal{O}(F\,d)$ encoder projection. Most prior optimization efforts focus on the decoder side. For example, Gao et al. (2025) introduce a fused sparse–dense TopK kernel that reduces wall-clock time and memory traffic, while Rajamanoharan et al. (2024a) decouple activation selection from magnitude prediction to improve the $\ell_{0}$–MSE trade-off. Separately, Rajamanoharan et al. (2024b) propose JumpReLU, a nonlinearity designed to mitigate shrinkage of large activations in sparse decoders.
Figure 1: Isoflops Pareto fronts for KronSAE vs. TopK SAE on Qwen-1.5B for different dictionary sizes $F$. 100 M tokens: KronSAE improves explained variance while heavily reducing parameters count. 500 M tokens: maintains gains across most sizes. 1000 M tokens: matches baseline quality with significantly fewer parameters; Stars denote TopK baselines and markers denote KronSAE configurations $(m,n,h)$. See details in section 4.1.
Conditional Computation.
Conditional-computation schemes avoid instantiating a full dictionary per token by routing inputs to a subset of expert SAEs via a lightweight gating network (Mudide et al., 2025) by employing the Mixture-of-Experts ideas (Shazeer et al., 2017), but still incur a dense per-expert encoder projection, leaving the encoder as the primary bottleneck.
Weight Factorization and Logical Activation Functions.
Outside of sparse decoders and Mixture-of-Experts, tensor-factorization methods have been used to compress large weight matrices in language models (Edalati et al., 2021). In parallel, differentiable-logic activations were introduced to approximate Boolean operators in a smooth manner (Lowe et al., 2021). Our method synthesizes these lines of work: we embed a differentiable AND-like gate into a Kronecker-factorized encoder to build rich, compositional features while preserving end-to-end differentiability.
3 Method
3.1 Preliminaries
Let $\mathbf{x}\in\mathbb{R}^{d}$ denote an activation vector drawn from a pretrained transformer. A conventional $\operatorname{TopK}$ SAE (Gao et al., 2025) produces a reconstruction $\hat{\mathbf{x}}$ of $\mathbf{x}$ via
$\mathbf{f}=\operatorname{TopK}\bigl{(}W_{\text{enc}}\mathbf{x}+\mathbf{b_{enc}}\bigr{)},\qquad\hat{\mathbf{x}}=W_{\text{dec}}\mathbf{f}+\mathbf{b_{dec}},$ (1)
where $W_{\text{enc}}\in\mathbb{R}^{F\times d}$ and $W_{\text{dec}}\in\mathbb{R}^{d\times F}$ are full-rank matrices and $\mathbf{f}\in\mathbb{R}^{F}$ retains the $k$ largest activations. The encoder cost therefore scales as $\mathcal{O}(Fd)$ per token.
3.2 KronSAE
In contrast to prior work that (i) accelerates decoding or (ii) sparsifies computation through routing, KronSAE directly reduces the encoder’s computational complexity by decomposing the $W_{\text{enc}}$ matrix into head-wise Kronecker factors and introducing a multiplicative $\operatorname{mAND}$ kernel. We decompose the latent space into $h$ independent components (e.g. heads). Each head $k$ is parameterised by the composition of two thin matrices $P^{k}\in\mathbb{R}^{m\times d}$ (composition base) and $Q^{k}\in\mathbb{R}^{n\times d}$ (composition extension), with dims $m<n\ll d$ and $F=h\,m\,n$. The pre-latents
$\mathbf{p}^{k}=\operatorname{ReLU}(P^{k}\mathbf{x}),\qquad\mathbf{q}^{k}=\operatorname{ReLU}(Q^{k}\mathbf{x})$ (2)
are combined through an element-wise interaction kernel independently in each head,
$\mathbf{z}^{k}_{i,j}:=\operatorname{mAND}(p^{k}_{i},q^{k}_{j}):=\begin{cases}\sqrt{p^{k}_{i}\,q^{k}_{j}},&p^{k}_{i}>0\text{ and }q^{k}_{j}>0,\\ 0,&\text{otherwise},\end{cases}$ (3)
where $\mathbf{z}^{k}\in\mathbb{R}^{m\times n}$. Flattening and concatenating across heads yields post-latents $\mathbf{f}\in\mathbb{R}^{F}$, to which the usual $\operatorname{TopK}$ operator is applied. The $\operatorname{mAND}$ kernel smoothly approximates Boolean AND gate, ensuring non-zero activation only if both inputs are positive, while preserving gradient flow and activation magnitude for stable reconstruction. Further insights into the behavior of $\operatorname{mAND}$ are provided in Appendix C.
The encoder cost per token drops from $\mathcal{O}(Fd)$ to $\mathcal{O}\bigl{(}h(m+n)d\bigr{)}$ (see Appendix A.2). KronSAE thus reduces FLOPs and parameter count without routing overhead, and is orthogonal to existing sparse decoder kernels (Gao et al., 2025) and thus can be combined with them for end-to-end speed-ups.
4 Experiments
Dictionary size m n h Activation Explained Variance
32768 2 4 4096 $\operatorname{mAND}(u,v)$ $\mathbf{0.83356}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.82669
$u\cdot v$ 0.82373
4 8 1024 $\operatorname{mAND}(u,v)$ $\mathbf{0.82197}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.81912
$u\cdot v$ 0.81432
65536 2 4 8192 $\operatorname{mAND}(u,v)$ $\mathbf{0.84445}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.83276
$u\cdot v$ 0.82974
4 8 2048 $\operatorname{mAND}(u,v)$ $\mathbf{0.83503}$
$\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$ 0.82968
$u\cdot v$ 0.8251
Table 1: Explained variance for KronSAE using different composition activations under a 1B-token training budget. Bold values indicate the best performance of $\operatorname{mAND}$ across composition kernels.
We train SAEs on the residual streams of Qwen-2.5-1.5B-Base (Yang et al., 2024), Pythia-1.4B-deduped (Biderman et al., 2023), and Gemma-2-2B (Team et al., 2024) language models. Activations are collected on FineWeb-Edu, a filtered subset of educational web pages from the FineWeb corpus (Penedo et al., 2024). We measure reconstruction quality via explained variance (EV),
$\mathrm{EV}=1-\frac{\operatorname{Var}(\mathbf{x}-\hat{\mathbf{x}})}{\operatorname{Var}(\mathbf{x})},$
so that $1.0$ is optimal, and use automated interpretability pipeline (Bills et al., 2023; Paulo et al., 2024) and SAE Bench (Karvonen et al., 2025) to evaluate properties of SAE features. Our experiments (see detailed setup in Appendices A and D) address three questions:
  1. 1.
    Does KronSAE yield higher EV than a TopK SAE under fixed compute?
  2. 2.
    Which design choices (nonlinearity, $(m,n,h)$) drive EV improvements?
  3. 3.
    How do these choices affect properties and interpretability of learned features?
4.1 Scaling Laws
We examine the scaling behavior of KronSAE against a conventional TopK SAE on Qwen-1.5B under equivalent FLOPs budgets. Models are grouped by dictionary size $F\in\{2^{15},2^{16},2^{17}\}$, and each KronSAE configuration $(m,n,h)$ matching $F$ hidden features is trained under token budgets adjusted to match the per-step FLOPs of training a TopK SAE for 100 M, 500 M, and 1000 M tokens (see FLOPs calculation details in Appendix A).
Figure 1 plots explained variance versus trainable parameter count across these isoflops budgets. Under a 100 M token budget, all KronSAE variants exceed the TopK baseline in reconstruction quality, achieving up to a 4.3% increase in explained variance with roughly 54.7% fewer parameters. At 500 M tokens, KronSAE maintains  0.8% advantage and 43.8% fewer parameters across most model sizes, and the compositional head structure keeps smaller configurations on or above the baseline Pareto front. With 1000 M tokens, although the TopK SAE narrows the gap, KronSAE still matches baseline reconstruction quality with approximately 46.1% fewer parameters. Smaller composition base dimension ($m=2$) show especially strong returns under tight compute constraints. We observed similar results on Pythia-1.4B (see section A.3).
4.2 Ablations
Figure 2: Iso-FLOPs ablation: EV against head count $h$ and base dimension $m$ for KronSAE under 100M, 500M, and 1B token budgets. Higher $h$ and smaller $m$ yield improved reconstruction quality.
Composition activation.
To isolate the impact of our $\operatorname{mAND}$ operator, we compare it to two simpler interaction kernels: (i) the element-wise product of ReLUs, $\operatorname{ReLU}(u)\cdot\operatorname{ReLU}(v)$, and (ii) the raw product $u\cdot v$. As reported in Table 1, under a 1B-token training budget and for both $(m,n,h)=(2,4,4096)$ and $(4,8,1024)$ configurations, the $\operatorname{mAND}$ variant consistently achieves the highest explained variance, outperforming the alternatives by a clear margin.
Decomposition hyperparameters.
Using the iso-FLOPs setup from Section 4.1, we systematically vary the number of heads $h$ and the per-head base dimension $m$ (with $n=F/(mh)$). Figure 2 shows that for large training budgets (500M–1B tokens), smaller $m$ (and thus larger $n$) yields higher reconstruction quality, since a smaller base dimension frees capacity for more expressive extension features. Under the tighter 100M-token regime, however, configurations with $m=4$ outperform those with $m=2$ or $m=8$, indicating a trade-off between per-head representational richness and data efficiency. Furthermore, fixing $m$ and increasing $h$ yields near-linear gains in explained variance. This corresponds to our intuition that each head acts as a component of the feature-space decomposition, efficiently storing a group of correlated features and allowing more fine-grained routing of activation patterns.
Sparsity and depth-wise position.
Finally, we assess KronSAE’s robustness to different sparsity levels and layer depths. Figure 3(a) compares KronSAE and TopK SAE across a range of $L_{0}$ sparsity settings on the 14th layer of Qwen-2.5-1.5B and the 12th layer of Gemma-2-2B; Figure 3(b) evaluates performance across layers in Qwen-2.5-1.5B. In every case, KronSAE matches or exceeds the reconstruction quality of the TopK baseline under the same FLOPs budget, demonstrating that our Kronecker-factorized encoder maintains its advantages regardless of sparsity level or depth.
4.3 Absorption
The primary challenge in interpretability is feature absorption, where one learned feature becomes a strict subset of another (e.g. a “Lion” feature is entirely subsumed by a “starts with L” feature) and consequently fails to activate on instances that satisfy the broader concept but not its superset representation (Chanin et al., 2024).
Figure 4 reports three absorption metrics across sparsity levels $\ell_{0}\in\{16,32,64,128,256\}$: (1) the mean absorption fraction, measuring the proportion of features that are partially absorbed; (2) the mean full-absorption score, quantifying complete subsumption events; and (3) the mean number of feature splits, indicating how often a single conceptual feature fragments into multiple activations. Across all $\ell_{0}$, KronSAE variants consistently reduce both the absorption fraction and the full-absorption score relative to the TopK SAE baseline, while maintaining a similar rate of feature splits. We attribute KronSAE’s improved disentanglement to two complementary design choices:
  1. 1.
    Smooth $\operatorname{mAND}$ activation. By emitting nonzero output only when both pre-latents are positive, we introduce a differentiable AND gate that prevents a broadly polysemantic primitive from entirely subsuming a more specific one. Consequently, composite post-latents fire mainly in the intersection of their constituent concepts, encouraging each pre-latent to specialize on a single semantic mode rather than inherit its “parent” activation region.
  2. 2.
    Head-wise Cartesian decomposition. Dividing the latent space into $h$ independent subspaces (each with its own $m\times n$ grid of primitive interactions) ensures that specialized concepts (such as “elephant”) are confined to a single head and cannot fully absorb more general concepts (such as “starts with E”) in another.
Figure 3: (a) EV versus sparsity level $L_{0}$ for KronSAE and TopK SAE on the 14th layer of Qwen-2.5-1.5B and the 12th layer of Gemma-2-2B under iso-FLOPs constraints. (b) EV across layers of Qwen-2.5-1.5B, demonstrating that KronSAE matches TopK performance regardless of depth.
Together, these mechanisms produce more monosemantic features, simplifying downstream causal interventions and targeted probing. Notably, the mean number of feature splits remains comparable to the TopK baseline, as Cartesian decomposition alone does not inherently alter the fragmentation of individual primitives (see Section 5.2).
Figure 4: Feature absorption metrics on Qwen-2.5-1.5B. KronSAE configurations (various $m,n$) exhibit lower mean absorption fractions and full-absorption scores across different $\ell_{0}$.
5 Analysis
In this section we investigate the properties of our architecture and of learned features, and compare it to TopK architecture. We show that pre-latents interactions are closely resemble the logic AND gate.
5.1 Toy Model of Superposition
Figure 5: We generate data with covariance matrix that consist of blocks with different sizes on diagonal and off diagonal (left panel). We then examine the decoder-weight covariance $W_{dec}\cdot W_{dec}^{T}$ to assess feature-embedding correlations. Panels (middle panels) show that a TopK SAE recovers these correlation structures only weakly, even after optimal atom matching. In contrast, KronSAE (right panel) more accurately reveals the original block patterns.
To evaluate how well different sparse autoencoder architectures recover known correlation patterns, we construct a controlled experiment using a synthetic, block-structured covariance model. Input vectors $\mathbf{x}_{\mathrm{sparse}}\in\mathbb{R}^{256}$ sampled under a heavy-tailed Bernoulli distribution with probability $p=0.875$ that value will be zero. We perform a Cholesky decomposition $S=LL^{\top}$ on the covariance matrix $S$ and set $\bar{\mathbf{x}}_{\mathrm{sparse}}=L\,\mathbf{x}_{\mathrm{sparse}},$ so that $\bar{\mathbf{x}}_{\mathrm{sparse}}$ exhibits the desired structure.
We train autoencoder (AE) to reconstruct $\mathbf{x}_{\mathrm{sparse}}$. For this we use small toy model following the (Elhage et al., 2022):
$\mathbf{\hat{x}}=\operatorname{ReLU}(W^{T}W\cdot\bar{\mathbf{x}}_{\mathrm{sparse}}+\mathbf{b}).$ (4)
We collect hidden states ($W\cdot\bar{\mathbf{x}}_{\mathrm{sparse}}$) from AE and then train TopK SAE and our proposed KronSAE with $|F|=256$ and $\text{topk}=4$ to reconstruct it. After training, we extract the decoder weight matrices $W_{\mathrm{dec}}$ from each model and compute the covariance $C_{\mathrm{dec}}=W_{\mathrm{dec}}W_{\mathrm{dec}}^{\top}$. To compare TopK SAE embeddings to the AE reference, we match atoms by minimal Euclidean distance, ensuring fair alignment before analysis. Result is shown in Figure 5.
[…]
5.2 Correlations in SAEs Trained on Language
To examine the correlation structure of features learned in our SAE, we have calculated the correlations on 5k texts from the training dataset. For each feature we calculate the mean correlation with features within its group (head) and with all other features, and compare the randomly initialized KronSAE with $m=4,n=4$ with the trained one. It turns out, as shown in Figure 6, that correlations are indeed significantly higher within a single head and higher than for random SAE, which suggest that our choice to impose the correlated structure in SAE latents works as intended.
Figure 6: Correlations between features in KronSAE with $m=4,n=4$ within a head and with features from other heads. Our design induces higher correlations within a group, which also gets stronger after training, although SAE have also learned correlated features from different heads.
5.3 Analysis of Learned Features
We have compared our KronSAE and TopK architecture in terms of interpretability and feature properties, and we have analyzed the properties of groups in KronSAE. For this, we choose the 14th layer of Qwen2.5-1.5B and a dictionary size of 32k features, of which the first 3072 were selected. KronSAE was chosen with $m=4,n=4$. We run for 24M tokens total to collect data. Our interpretation pipeline follows the common methodology: LLM interprets the activation patterns (Bills et al., 2023) and we evaluate obtained interpretations using the detection score and the fuzzing score Paulo et al. (2024).
For each selected feature, among the standard mean activation value and frequency, we calculate two additional metrics. Low values of token entropy suggest that feature activates more frequently on small number of tokens, thus it is token-specific; high value of multitoken ratio indicates that feature tends to activate multiple times in a single sentence. We have observed that both these metrics have non-negligible negative correlation with the final interpretability scores.
For more details on the data collection and interpretation pipeline, see Appendix D. For additional analysis of properties of learned features, see Appendix E.
SAE properties and encoding mechanism.
We observe that the features learned by KronSAE are more specific, indicated by lower values of the computed metrics and higher interpretability scores, as shown in Figure 7. Since post-latents are significantly more interpretable than corresponding pre-latents, we hypothesize the hidden mechanism for encoding and retrieval of the required semantics.
Figure 7: Distribution of properties for TopK SAE and KronSAE ($m=4,n=4$) with 32k dictionary size trained on Qwen2.5-1.5B. Pre and Post suffixes denote pre- and post-Cartesian product latents. Our SAE achieves better interpretability scores by learning specialized feature groups, indicated by lower activation frequency and lower variance in activated tokens.
By examining activating examples and interpretations of latents, we observe that pre-latents may carry multiple distinct and identifiable modes of activation, such as composition base element 3 in head 23 shown in Table 2, and be very abstract compared to resulting post-latents. Polysemanticity of pre-latents is expected to be a consequence of reduced "working" number of encoder latents, since we decompose the full dictionary size and reduce the encoder capacity.
Thus, we hypothesize that the encoding of specific semantics in our SAE may be done via magnitude, which we validate by examining the activation examples. For the above mentioned pre-latent, the "comparison" part is encoded in the top 75% quantile, while the "spiritual" part is mostly met in the top 25% quantile, and the "geographical" part is mainly encoded in the interquartile range. We also consider but do not investigate the possibility that it may depend on the context, e.g. when the model uses the same linear direction to encode different concepts when different texts are passed to it.
Semantic retrieval and interpretable interactions.
Heads usually contain a groups of semantically related pre-latents, e.g. in head 136 there are three base elements and one extension covering numbers and ordinality, two extension elements related to geographical and spatial matters, one question-related base and one growth-related extension. Interestingly, most post-latents for this head have higher interpretability score than both its parent pre-latents, which is unusual.
The retrieval happens primarily via the mechanism closely resembling the logical AND circuit, where some pre-latent works as the bearer of multiple semantics, and the corresponding pre-latent (base or extension) works as specifier. An illustrative example is shown in Table 2: we see that the base contains three detectable sub-semantics, and each extension then retrieves the particular semantics.
Component Interpretation Score
Base 3 Suffix “-like” for comparative descriptors, directional terms indicating geographical regions, and concepts related to spiritual or metaphysical dimensions 0.84
Extension elements and their compositions with base 3
Extension 0 Interpretation: Comparative expressions involving “than” and “as” in contrastive or proportional relationships. 0.87
Composition: Similarity or analogy through the suffix “-like” across diverse contexts. 0.89
Extension 1 Interpretation: Specific terms, names, and abbreviations that are contextually salient and uniquely identifiable. 0.66
Composition: Medical terminology related to steroids, hormones, and their derivatives. 0.84
Extension 2 Interpretation: Spiritual concepts and the conjunction “as” in varied syntactic roles. 0.80
Composition: Abstract concepts tied to spirituality, consciousness, and metaphysical essence. 0.93
Extension 3 Interpretation: Directional and regional descriptors indicating geographical locations or cultural contexts. 0.84
Composition: Directional terms indicating geographical or regional divisions. 0.91
Table 2: Interactions between composition base element 3 in head 23 and all extension elements in that head. Interaction happens in a way that closely resembles the Boolean AND operation: base pre-latent is polysemous, and the composition post-latent is the intersection, i.e. logical AND between parent pre-latents. See details in Section 5.3.
Other types of interaction may occur, such as appearance of completely new semantics, for example composition between base 3 and extension 1 in Table 2 where medical terminology arises and could not be interpreted as simple intersection between two pre-latents semantics. Another example is a case of head 3 where base 3 has sub-semantics related to technical instruments and extension 2 have semantics related to the posession and necessity, and their combination gives the therapy and treatment semantics which looks more like addition than intersection.
It is a frequent case that post-latent inherit semantics of only one parent, or the impact of another parent is not detectable, which usually happens if parent has a very broad interpretation and low score. However, it requires more sophisticated techniques to properly identify the fine-grained structure of interactions than just looking at the resulting latent descriptions, so we leave it to further work. Despite this, the AND-like gate is a very common behavior. See more examples in Appendix G.
Geometry of post-latents.
Each post-latent vector has a vector representation in the residual stream represented by the corresponding column in $W_{\text{dec}}$, which is the approximation of overcomplete basis vectors we search for when training SAEs. We had not observed any notable differences in feature geometry between TopK and our SAEs, except that our architectural design leads to clustering so that post-latents produced by same head, base or a extension elements are grouped in a tight cluster, and the geometry is dependent on hyperparameters $h,m,n$ we choose, which is expected and may be useful for further applications such as steering. See more details in Appendix E.
6 Conclusion and Future Work
We introduce KronSAE, a sparse autoencoder architecture design that directly tackles the long-standing encoder inefficiency bottleneck by combining head-wise Kronecker factorisation with a $\operatorname{mAND}$ gating. This significantly lowers number of parameters compared to standard TopK SAEs, while improving reconstruction fidelity and yielding more interpretable features by utilizing their correlations. Our analysis links these gains to the complementary effects of compositional latent structure and logical AND-style interactions, offering a new lens on how sparsity and factorisation can synergise in representation learning.
Limitations.
Despite its benefits, KronSAE has several limitations. Its gains depend on careful tuning of $(m,n,h)$ and the mAND activation—misconfigured settings can negate both efficiency and quality improvements. Our evaluation is limited to mid-sized transformer models and a single web corpus, so its applicability to larger models, other domains, or languages remains untested.
Future Work.
We identify three directions for extending this work: (i) Transcoding. Treat transcoders (Dunefsky et al., 2024) as implicit routers of information and investigate alternative logical gating functions (e.g. XOR or composite gates) to improve interpretability and circuit analysis. (ii) Crosscoding. Generalize KronSAE to a crosscoder setting (Lindsey et al., 2024) uncover interpretable, cross-level compositionality via logic operations. (iii) Dynamic Composition. Explore learnable tuning of both the number of attention heads and their dimensionality, enabling fine-grained decomposition into groups of correlated features at varying scales.
Appendix A More Details About Experiments
A.1 Experimental setup.
Training details.
All SAEs are optimized using AdamW with an initial learning rate of $8\times 10^{-4}$, a cosine learning-rate schedule with a minimum LR of $1\times 10^{-6}$, and a linear warmup for the first 10% of total training steps. We use a global batch size of 8,192. We sweep over dictionary (latent) sizes of $2^{15}$, $2^{16}$, and $2^{17}$ features. For our KronSAE variant, we further sweep the number of heads $h$ and the per-head dimensions $m$ and $n$ such that $h\cdot m\cdot n$ equals the desired dictionary size. Regularization weights and auxiliary loss coefficients are kept constant throughout the runs to isolate the impact of architectural choices.
For all experiments, we spent about 330 GPU days on NVIDIA H100 80GB GPUs, including preliminary research.
SAE.
For all experiments on Qwen-2.5, we train each SAE on activations from layer 14. Also for Pythia-1.4B we use layer 14 and for Gemma-2-2B we take activations from layer 12. For most of our experiments, we use sparsity level of $\ell_{0}=50$ non-zero activations per token.
Initialization.
As observed by Gao et al. [2025], initializing the decoder as the transpose of the encoder ($W_{\mathrm{dec}}=W_{\mathrm{enc}}^{\top}$) provides a strong metric improvement. We adopt this strategy within KronSAE by partitioning $W_{\mathrm{enc}}$ into $h$ head-wise blocks of shapes $m\times d$ and $n\times d$, denoted $\{P_{i},Q_{i}\}_{i=1}^{h}$. For each head $k$, we define its decoded rows via a simple additive composition:
$C_{k}[i,j]\;=\;P_{k,i}\;+\;Q_{k,j},\quad i=1,\dots,m,\;j=1,\dots,n.$
Finally, flattening the matrices $\{C_{k}\}$ yields full decoder weight matrix $W_{\mathrm{dec}}\in\mathbb{R}^{F\times d}$.
A.2 FLOPs Calculation.
For TopK SAE and KronSAE we compute FLOPs in the following way:
$\operatorname{FLOPS_{TopK}}(d,F,k)$ $=dF+kd,$ (5)
$\operatorname{FLOPS_{KronSAE}}(d,m,n,h,k)$ $=dh(m+n)+mnh+kd\approx dh(m+n)+kd.$
We calculate flops for most effective variant of TopK where we perform vector matrix multipication only for nonzero activations.
A.3 Pythia model.
For Pythia-1.4B we train all SAEs on the 12th transformer layer. Figure 8 shows that, under larger equivalent token budgets (500M and 1000M), KronSAE achieves a wider performance margin over the TopK SAE than we observed with Qwen-2.5.
Figure 8: Isoflops comparison of KronSAE and TopK SAE on Pythia-1.4B across three token budgets.
Appendix B More Results on Synthetic
We further evaluate KronSAE on several variant block-diagonal covariance matrices (Figure 9). In each case, the decoder-weight covariance $C_{\mathrm{dec}}=W_{\mathrm{dec}}W_{\mathrm{dec}}^{\top}$ of KronSAE more faithfully reproduces the ground truth groupings than the TopK SAE. Notably, on the third covariance pattern (where some blocks are very small) TopK’s learned correlations nearly vanish, whereas KronSAE still uncovers the correct block structure. For the first covariance matrix, KronSAE yields sharply elevated correlations in the regions corresponding to the true blocks, in line with our design goal of head-wise compositionality.
Table 3 quantifies these observations via the RV coefficient and permutation tests. Even after optimally matching TopK’s atoms to a dense AE reference, TopK SAE attains only weak correlation alignment ($RV\approx 0.05-0.08$) with non-significant or marginal p-values. In contrast, KronSAE configurations achieve $RV$ values between 0.11 and 0.35 (all $p<0.001$), representing a 3–6x improvement in correlation recovery. These results confirm that our compositional encoder not only accelerates training but also robustly captures the intended hierarchical feature interactions across diverse covariance regimes.
Figure 9: Comparision of how different KronSAE try to reveal hidden structure of defined covariance matrix. The KronSAE models recover the underlying block-structured correlations more faithfully than the TopK baseline, with the finer head/composition split $(h=4,m=2)$ capturing smaller feature groups more accurately.
Covariance matrix SAE $RV$ p-value
1 TopK $0.046$ $0.1324$
TopK Matched $0.051$ $0.0102$
KronSAE $h=2,m=4,n=32$ $0.200$ $0.0002$
KronSAE $h=4,m=2,n=32$ $0.150$ $0.0002$
2 TopK $0.033$ $0.2212$
TopK Matched $0.080$ $0.0002$
KronSAE $h=2,m=4,n=32$ $0.221$ $0.0002$
KronSAE $h=4,m=2,n=32$ $0.319$ $0.0002$
3 TopK $0.035$ $0.2490$
TopK Matched $0.043$ $0.0002$
KronSAE $h=2,m=4,n=32$ $0.111$ $0.0002$
KronSAE $h=4,m=2,n=32$ $0.212$ $0.0002$
4 TopK $0.034$ $0.4613$
TopK Matched $0.043$ $0.0002$
KronSAE $h=2,m=4,n=32$ $0.346$ $0.0002$
KronSAE $h=4,m=2,n=32$ $0.334$ $0.0002$
Table 3: RV coefficient and permutation p-values for correlation recovery on four synthetic covariance patterns. KronSAE outperforms both the standard and atom-matched TopK SAE by a large margin, achieving statistically significant alignment ($p<10^{-3}$) across all cases.
Appendix C mAND as a Logical Operator
For KronSAE, we replace the original $\operatorname{AND_{AIL}}$  [Lowe et al., 2021] with a more restrictive approximation, $\operatorname{mAND}$. Since our objective is to drive each atom toward a distinct, monosemantic feature, we found that tightening the logical conjunction encourages sharper feature separation. Moreover, by using the geometric mean ($\sqrt{p\,q}$) rather than a simple product or minimum, $\operatorname{mAND}$ preserves activation magnitudes. A visual comparison of $\operatorname{mAND}$ and $\operatorname{AND_{AIL}}$ appears in Figure 10.
Figure 10: Comparison of the smooth $\operatorname{mAND}$ operator against the $\operatorname{AND_{AIL}}$  [Lowe et al., 2021].
Appendix D Feature Analysis Methodology
We analyze learned features using an established pipeline Bills et al. [2023], Paulo et al. [2024] consisting of three stages: (1) statistical property collection, (2) automatic activation pattern interpretation, and (3) interpretation evaluation. The following subsections detail our implementation.
D.1 Data collection
Our collection process uses a fixed-size buffer $B=384$ per feature, continuing until processing a predetermined maximum token count $T_{max}$. The procedure operates as follows:
Initial processing batches generate large activation packs of 1M examples, where each example comprises 256-token text segments. When encountering feature activations, we add them to the buffer, applying random downsampling to maintain size $B$ when exceeding capacity. This approach enables processing arbitrary token volumes while handling rare features that may require extensive sampling.
During collection, we compute online statistics including activation minimums, maximums, means, and frequencies. Post-processing yields two key metrics: token entropy and multitoken ratio. The token entropy is calculated as:
$\text{token entropy}=-\sum_{i=0}^{n}p_{i}\cdot\log(p_{i}),\quad p_{i}=\frac{\text{activations of token i}}{\text{total amount of activations}},$ (6)
where $n$ represents unique activated tokens. The multitoken ratio is:
multitoken $\text{multitoken ratio}=\frac{1}{b}\sum_{i=0}^{b}\frac{\text{number of activations in sequence $i$}}{\text{total tokens in sequence $i$}},$ (7)
with $b<B$ denoting collected context examples per feature.
We then segment examples using a 31-token context window (15 tokens before/after each activation), potentially creating overlapping but non-duplicated examples. Features with high multitoken ratio may have number of examples significantly exceeding $B$.
A separate negative examples buffer captures non-activating contexts. Future enhancements could employ predictive modeling (e.g., using frequent active tokens) to strategically populate this buffer with expected-but-inactive contexts, potentially improving interpretation quality.
D.2 Feature interpretations
For each feature, we generate interpretations by sampling 16 random activation examples above the median activation quantile and presenting them to Qwen3 14B [Yang et al., 2025] (AWQ-quantized with reasoning enabled). The model produces concise descriptions of the activation patterns. Empirical observations suggest reasoning mode improves interpretation quality, though we lack quantitative measurements. This aligns with findings in [Paulo et al., 2024], which compared standard one-sentence responses with Chain-of-Thought outputs, making model reasoning an interesting direction for future research.
The interpretation process uses the system prompt presented in a Figure 11. User prompts include all special characters verbatim, as some features activate specifically on these characters. A representative (slightly abbreviated) user prompt example is presented on Figure 12.
You are a meticulous AI researcher conducting an important investigation into patterns found in language. Your task is to analyze text and provide an explanation that thoroughly encapsulates possible patterns found in it.
Guidelines:
You will be given a list of text examples on which special words are selected and between delimiters like <<this>>. If a sequence of consecutive tokens all are important, the entire sequence of tokens will be contained between delimiters <<just like this>>. How important each token is for the behavior is listed after each example in parentheses.
- Your explanation should be a concise STANDALONE PHRASE that describes observed patterns.
- Focus on the essence of what patterns, concepts and contexts are present in the examples.
- Do NOT mention the texts, examples, activations or the feature itself in your explanation.
- Do NOT write "these texts", "feature detects", "the patterns suggest", "activates" or something like that.
- Do not write what the feature does, e.g. instead of "detects heart diseases in medical reports" write "heart diseases in medical reports".
- Write explanation in the last line exactly after the [EXPLANATION]:
Figure 11: System prompt for feature interpretations.
Examples of activations:
Text: ’ Leno<<,>> a San Francisco Democrat<<, said in a statement.>>’
Activations: ’ said (22.74), statement (27.84), in (27.54)’
Text: ’ city spokesman Tyler Gamble<< said in an>> email.’
Activations: ’ said (2.92), in (12.81), an (14.91)’
Text: ’ towpath at Brentford Lock. <<Speaking>> on BBC London 94’
Activations: ’Speaking (3.48)’
Text: ’ Michelle, a quadriplegic,<< told>> DrBicuspid.com’
Activations: ’ told (4.05)’
Text: ’ CEO Yves Carcelle<< said in a statement>>.’
Activations: ’ said (19.64), in (29.09), statement (29.39)’
Figure 12: Example of user prompt passed to LLM. This feature with 16 examples received the interpretation "Structural elements in discourse, including speech attribution, prepositional phrases, and formal contextual markers" with a detection score of 0.84 and fuzzing score of 0.76.
D.3 Evaluation pipeline
We evaluate interpretations using balanced sets of up to 64 positive (activation quantile > 0.5) and 64 negative examples, employing the same model without reasoning to reduce computational costs. When insufficient examples exist, we maintain class balance by equalizing positive and negative counts. The evaluation uses modified system prompts from [Paulo et al., 2024], with added emphasis on returning Python lists matching the input example count exactly. We discard entire batches if responses are unparseable or contain fewer labels than the number of provided examples.
We calculate two scores.
Detection Score: After shuffling positive/negative examples, we present up to 8 unformatted text examples per batch to the model. The model predicts activations (1/0) for each example, generating up to 128 true/predicted label pairs. The score calculates as:
score=12(correctly predicted positivestotal positives+correctly predicted negativestotal $\text{score}=\frac{1}{2}\left(\frac{\text{correctly predicted positives}}{\text{total positives}}+\frac{\text{correctly predicted negatives}}{\text{total negatives}}\right).$ (8)
Fuzzing Score: We <<highlight>> activated tokens on sampled examples, from which 50% are correctly labeled positive examples, 25% are mislabeled positive examples, and 25% are randomly labeled negative examples. We present batches of up to 8 examples and the model identifies correct/incorrect labeling, with scoring following Equation 8.
Appendix E Additional Feature Analysis Results
Feature property correlations.
Our analysis reveals significant correlations between feature properties and interpretability scores (Figure 13). Notably, token entropy and mean activation show substantial correlations with interpretability scores, suggesting their potential as proxies for assessing feature quality without running the full interpretation pipeline. These findings are based on analysis of the first 3072 features from 32k TopK and KronSAE (m=4, n=4) trained on 24M tokens, warranting further validation with larger-scale studies.
Pre-latent to post-latent relationships.
We investigate how post-latent properties correlate with various combinations of pre-latent properties, including individual values, means, products, and the $\operatorname{mAND}$ operation (product followed by square root). Figure 14 demonstrates that post-latent multitoken ratio, token entropy, and frequency show stronger correlations with pre-latent products or $\operatorname{mAND}$ values than with individual pre-latent properties or their means.
Figure 13: Correlation coefficients (Pearson and Spearman) between properties of TopK and KronSAE latents. Token entropy emerges as a strong predictor of interpretability scores, while higher mean activation and lower frequency also indicate more interpretable features.
Figure 14: Correlation patterns between properties of post-latents and pre-latents.
Basis geometry.
As noted in Section 5.3, latent embeddings primarily exhibit clustering within their originating groups (head, base, extension). With the support of observations reported in Sections 4.2 and 5.3, we find that models with more heads achieve better reconstruction while producing more diverse basis vectors. This suggests that fine-grained architectures yield more expressive representations, although they may also exhibit undesired challenging behavior like feature splitting [Bricken et al., 2023] or absorption [Chanin et al., 2024].
Figure 15 visualizes this structure through UMAP projections (n_neighbors=15, min_dist=0.05, metric=’cosine’) of decoder weights from the first 8 heads of 32k SAEs with varying m,n configurations. The plots reveal distinct clustering patterns: for $m<n$ we observe tight base-wise clustering with weaker grouping by extension, and for $m\geq n$ extension-wise clustering is stronger.
Figure 15: UMAP visualization of post-latent clustering patterns by head, base, and extension group membership. We observe tight clusters by base for $m<n$ and by extension for $m\geq n$.
This asymmetry suggests that pre-latent capacity requirements directly manifest in the embedding geometry - components with lower polysemanticity (extensions when m < n) exhibit greater geometric diversity. We expect symmetric behavior for reciprocal configurations (e.g., m=4,n=8 vs. m=8,n=4), merely swapping the roles of bases and extensions.
Appendix F KronSAE in Terms of Tensor Diagram
The proposed encoder architecture can be visualized as a tensor diagram (Figure 16). Notably, this formulation draws a connection to quantum mechanics, where |$\ket{\mathbf{f}}$ represents the (unnormalized) state of two disentangled qubits described by |$\ket{\boldsymbol{p}}$ and |$\ket{\boldsymbol{q}}$.
If we were to sum the outputs of the encoder’s heads instead of concatenating them, |$\ket{\mathbf{f}}$ would correspond to a separable quantum state. This scenario can be expressed via the Schmidt decomposition:
|$\ket{\mathbf{f}}=\sum_{h}\ket{\boldsymbol{p}_{h}}\otimes_{K}\ket{\boldsymbol{q}_{h}},$
where $\otimes_{K}$ denotes the Kronecker product. However, preliminary experiments revealed that this alternative design results in poorer performance compared to the concatenation-based approach.
Figure 16: For a single head, the KronSAE encoder architecture separates the input $\boldsymbol{x}$ into two distinct components, $\boldsymbol{p}$ and $\boldsymbol{q}$, via matrix multiplications with $W_{enc}^{p}$ and $W_{enc}^{q}$ accordingly. These components are then combined via the outer product $\boldsymbol{p}\otimes\boldsymbol{q}$, resulting in a matrix representation. To produce the final output vector $\mathbf{f}$, this matrix is flattened into a single vector using a multi-index mapping.
Appendix G Analysis of Compositional Structure
Here we analyze more examples of interactions in various heads.
Head 3.
For this head we have selected all base elements and extension 2, shown in Table 4. Extension element 2 shows moderate interpretability with clear AND-like interactions: with base 1 (semantic inheritance through shared pre-latent semantics) and base 2 (retaining only instrument-related semantics). Notable interactions occur with base 0 (acquiring medical semantics while preserving metric/number aspects) and base 3 (combining instrument semantics with necessity to yield therapy/treatment concepts). The high interpretability scores suggest potential additional encoding mechanisms beyond simple intersection, possibly related to activation magnitude, though dataset or interpretation artifacts cannot be ruled out without further validation.
Component Interpretation Score
Extension 2 Scientific instruments, acronyms, and critical numerical values in technical and astronomical contexts 0.71
Base elements and their compositions with extension 2
Base 0 Interpretation: Punctuation marks and line breaks serving as structural separators in text. 0.66
Composition: Health-related metrics focusing on survival rates, life expectancy, and longevity. 0.88
Base 1 Interpretation: Numerical values, both in digit form and as spelled-out numbers, often accompanied by punctuation like decimals or commas, in contexts of measurements, statistics, or quantitative expressions. 0.80
Composition: Numerical digits and decimal points within quantitative values. 0.86
Base 2 Interpretation: Nuanced actions and adverbial emphasis in descriptive contexts. 0.71
Composition: Astronomical instruments and their components, such as space telescopes and their acronyms, in scientific and observational contexts. 0.90
Base 3 Interpretation: Forms of the verb "to have" indicating possession, necessity, or occurrence in diverse contexts. 0.91
Composition: Antiretroviral therapy components, viral infection terms, and medical treatment terminology. 0.87
Table 4: Interactions between extension 2 in head 3 and all base elements in that head.
Head 136.
This head exhibits higher interpretability in post-latents than pre-latents. Key observations from the Table 5 include: extension 2 with base 0 narrows semantics to Illinois (likely inheriting geographical subsemantics), while interactions with bases 2-3 demonstrate complexity beyond simple intersection, often introducing additional semantics requiring deeper investigation.
Component Interpretation Score
Extension 2 Hierarchical scopes, geographic references, and spatial dispersal terms 0.78
Base elements and their compositions with extension 2
Base 0 Interpretation: Numerical decimal digits in quantitative expressions and proper nouns. 0.79
Composition: The state of Illinois in diverse contexts with high significance. 0.95
Base 1 Interpretation: The number three and its various representations, including digits, Roman numerals, and related linguistic forms. 0.84
Composition: Geographic place names and their linguistic variations in textual contexts. 0.91
Base 2 Interpretation: Ordinal suffixes and temporal markers in historical or chronological contexts. 0.87
Composition: Terms indicating layers, degrees, or contexts of existence or operation across scientific, organizational, and conceptual domains. 0.82
Base 3 Interpretation: Question formats and topic introductions with specific terms like "What", "is", "of", "the", "Types", "About" in structured text segments. 0.77
Composition: Spatial spread and occurrence of species or phenomena across environments. 0.87
Table 5: Interactions between extension 2 in head 136 and all base elements in that head.
Head 177.
Latents presented in Table 6 emonstrates more consistent AND-like behavior than Heads 3 and 136, closely matching the interaction pattern shown in Figure 2, supporting our theoretical framework.
Component Interpretation Score
Extension 1 Geographical mapping terminology and institutional names, phrases involving spatial representation and academic/organizational contexts 0.90
Base elements and their compositions with extension 1
Base 0 Interpretation: Proper nouns, abbreviations, and specific named entities. 0.64
Composition: Geographical or spatial references using the term "map". 0.93
Base 1 Interpretation: Emphasis on terms indicating feasibility and organizations. 0.80
Composition: Specific organizations and societal contexts. 0.89
Base 2 Interpretation: Institutional names and academic organizations, particularly those containing "Institute" or its abbreviations, often paired with prepositions like "of" or "for" to denote specialization or affiliation. 0.89
Composition: Institutional names containing "Institute" as a core term, often followed by prepositions or additional descriptors. 0.92
Base 3 Interpretation: Closure and termination processes, initiating actions. 0.79
Composition: Initiating or establishing a state, direction, or foundation through action. 0.85
Table 6: Interactions between extension 1 in head 177 and all base elements in that head.